OpenStudy (anonymous):

If a computer finishes a job in 2 hours and another computer finishes a job in 5 hours how long will it take for them to finish together?

OpenStudy (anonymous):

Are you sure? That seems really simple to do. I thought it would be more complicated :/

OpenStudy (anonymous):

Are you sure @monokerous

OpenStudy (anonymous):

it's not right 2 people will always be more efficient than 1

OpenStudy (anonymous):

Oops! Wrong interpretation. Sorry, sorry. :/

OpenStudy (anonymous):

Can someone help me please? :(

OpenStudy (anonymous):

Computer's A one hour work: \(\frac{1}{2}\) Computer's B one hour work : \(\frac{1}{5}\) {A + B)'s 1 hour work : \(\frac{1}{2} + \frac{1}{5} \implies \frac{7}{10}\) So both will complete in : \(\frac{10}{7}\) hours.

OpenStudy (anonymous):

Would i reduce that to 1 3/7 hours or just keep it as it is?

OpenStudy (anonymous):

Yes you can do that..

OpenStudy (anonymous):

Ahhh okay, thank you! :)

OpenStudy (anonymous):

Welcome dear..